OpenStudy (anonymous):

What is the value of y? 6(5 + 3) = 6 • 5 + 6 • y y=

OpenStudy (arindameducationusc):

you there?

OpenStudy (arindameducationusc):

48=30+6y

OpenStudy (arindameducationusc):

18/6=3

OpenStudy (arindameducationusc):

3 is the answer

OpenStudy (imtiaz7):

6(5 + 3) = 6 • 5 + 6 • y 30+18=30+6.y 48-30=6y 18=6y y=18/6 y=3

OpenStudy (anonymous):

thanks both of you :)